OpenStudy (tgstudios):

The sides of a square are 2 4/9 inches long. What is the area of the square? 2 16/81 sq in 4 16/81 sq in 2 8/9 sq in 4 8/9 sq in Please Help!!

OpenStudy (seratul):

The formula for area of a square is Side^4 So basically do 2 4/9 * 2 4/9 * 2 4/9 * 2 4/9

OpenStudy (tgstudios):

the answer choices?

OpenStudy (seratul):

Did you do the math? 2 4/9 * 2 4/9 * 2 4/9 * 2 4/9 What did you get?

OpenStudy (tgstudios):

2 4/9 times itself 4 times is 35.something... the answer coices are above, that's not even close

OpenStudy (seratul):

Whoops, i meant to say 2 4/9 * 2 4/9

OpenStudy (tgstudios):

I did that too, it gives me 5.975308...

OpenStudy (tgstudios):

just putting the answer choices closer: 2 16/81 sq in 4 16/81 sq in 2 8/9 sq in 4 8/9 sq in

OpenStudy (seratul):

You shouldve gotten 4.888

OpenStudy (tgstudios):

I didnt

OpenStudy (tgstudios):

2 4/9 * 2 4/9 = 5.975

OpenStudy (seratul):

You did the math wrong then. 2 4/9 is equal to 22/9 which is equal to 2.444 2.444 times 2.444 is 4.888

OpenStudy (tgstudios):

that's addition dude... XD

OpenStudy (seratul):

Oh wow. I'm dumb. I'm extremely sorry.

OpenStudy (tgstudios):

heres the problem, my mom thinks the exponents are exponents, I don't think they are

OpenStudy (seratul):

That explains it, it is exponents.

OpenStudy (tgstudios):

but I did that too... XD XD

OpenStudy (seratul):

Did you put parenthesis around the exponents?

OpenStudy (tgstudios):

yup, gives me .750123457

OpenStudy (seratul):

Google also says that the area should be 5.97 There is something wrong about those answer choices.

OpenStudy (tgstudios):

XD this question is so messed up!

OpenStudy (seratul):

Oh, I see why lmao. 2^(4/9) is equal to 1.36079

OpenStudy (seratul):

If you square that, you get 1.8517

OpenStudy (tgstudios):

no its not...

OpenStudy (seratul):

Yes it is, I just did it on my calculator

OpenStudy (tgstudios):

4/9 is less than half, half of 2 is 1

OpenStudy (seratul):

Put it into the calculator, you will get the same answer as well.

OpenStudy (tgstudios):

just putting the answer choices closer: 2 16/81 sq in 4 16/81 sq in 2 8/9 sq in 4 8/9 sq in

OpenStudy (seratul):

I got the answer btw. Its C

OpenStudy (seratul):

Do you get why?

OpenStudy (tgstudios):

no, but ok, well ill tell you if your right! thanks!

OpenStudy (seratul):

Tell me what you get when you do 2^(8/9)

OpenStudy (seratul):

Alright man.

OpenStudy (tgstudios):

1.77777777777777 :)

OpenStudy (seratul):

And when you square it?

OpenStudy (seratul):

Wait, you multiplied. You didnt do it to the exponent.

OpenStudy (tgstudios):

3.1604

OpenStudy (seratul):

You multiplied. It should be an exponent.

OpenStudy (tgstudios):

1.857

OpenStudy (seratul):

Now square it

OpenStudy (tgstudios):

3.448448448448...

OpenStudy (seratul):

No, you did it wrong lol. 2^(4/9) is equal to 1.36079

OpenStudy (seratul):

when you square it, you get 1.851749425 which is equal to 2^(8/9)

Directrix (directrix):

I did 2 4/9 * 2 4/9 = 22/9 * 22/9 = 484/81 = 5 79/81 -------- I also did (2+(4/9))* (2+(4/9)) and got the same answer. I am thinking that the correct answer is not given. Whoever wrote the problem thought that the correct answer is 4 16/81 and got that answer by thinking squaring 2 4/9 means to square the 2 and then to square the 4/9. That is not correct.

OpenStudy (seratul):

the 4/9 is an exponent..

OpenStudy (seratul):

@Directrix

OpenStudy (seratul):

If you put it as an exponent, you get 1.36079 Then you have to square it cause it is a square. You get 1.85174925 This is equivalent to choice C which is 2^(8/9)

Directrix (directrix):

Why not use the properties of exponents?